OpenStudy (anonymous):

find the equation of the line tangent to cos(x+y) = 2x+2y at the point (pi,pi)

OpenStudy (anonymous):

Are you comfortable with implicit differentiation at all, or do we need help on that part?

OpenStudy (anonymous):

I need help with implicit differentiation

OpenStudy (anonymous):

Alright. Well, as always, we want to find dy/dx, or y' if you will. Except this time we don't have y by itself. Usually, if we have something like y = x^2, we get dy/dx = 2x and we're done. This time, though, we have y's scattered about and everytime we differentiate one of those y's we're going to get a dy/dx. Well, the idea is to differentiate everything, get all of those dy/dx's collected together and then move everything else to the other side, essentially treating dy/dx as a variable in which we're trying to solve for. So that's the plan. So to start out, the derivative of cos(x+y). I'll need to differentiate this with chain rule, taking the derivative of cosine first and then multiplying it by the derivative of x+y. Think you could differentiate cos(x+y)?

OpenStudy (anonymous):

so first we need to take the derivative of cos (x+y) ?

OpenStudy (anonymous):

I know the derivative of cos is -sinx

OpenStudy (anonymous):

We'll need to take the derivative of all of the terms in the problem, using any chain rule, product rule, or whatever necessary. While doing this, if we ever apply a derivative rule to a y variable, we'll need to tack on dy/dx.

OpenStudy (anonymous):

Right, -sinx. But we want to maintain the inner function. Whenever we take derivative, any inner function or inner portion doesn't get changed. So we will get -sin(x+y), the x+ y part unchanged. But now we must apply chain rule. x+y is an inner function, so we must find its derivative and multiply it by -sin(x+y). So what would the derivative of x+y be?

OpenStudy (anonymous):

(y' (x)+1)(-sin(x+y)

OpenStudy (anonymous):

Yep, that works. I'll be writing that as -sin(x+y)(1+y') if you don't mind, just to try and not have a ton of clutter. But exactly, you got it. So now the next two derivatives are simple. We know the derivative of 2x = 2 and the derivative of 2y =2, but since it was a y derivative, we attach y'. So all together we get: \[-\sin(x+y)(1+y') = 2x+2y'\] So now we have two y's in our equation. We have to solve for y' and get it all by itself. Think you can do that part?

OpenStudy (anonymous):

so am I supposed to factor the equation out?

OpenStudy (anonymous):

Yes. Just like if we had something like 2x - 5 = 3x + 2, we would want to gather all the x's and then get it by itself. We do the same thing with y', get them all factored out and on one side of the equation by itself.

OpenStudy (anonymous):

ok I understand your example but I get confused with the sin part

OpenStudy (anonymous):

Well, whatever is inside of the angle of sin cannot be touched. -sin(x+y) is all one unit, it cannot be broken apart without identities, and we don't want to do that. So when we distribute, we will have\[-\sin(x+y)(1+y') = -\sin(x+y) - y'\sin(x+y)\]

OpenStudy (anonymous):

ok how did you get -sin (x+y) on the right side?

OpenStudy (anonymous):

Well, I was only focused on the sin(x+y) portion of the problem since that was the part that seemed to confuse you. The whole equation I would have is: \[-\sin(x+y) -y'\sin(x+y) = 2x + 2y'\]

OpenStudy (anonymous):

ohh ok I see! that was only half of it thanks!

OpenStudy (anonymous):

Yeah, exactly :)

OpenStudy (anonymous):

Now we just need to isolate y' on one side of the equation

OpenStudy (anonymous):

ok i will do that

OpenStudy (anonymous):

Show me what ya got when you finish and we can continue :)

OpenStudy (anonymous):

ok so the first step to isolating y' would be to subtract 2x on both sides right? and then we can add y' on both sides to leave y' only on the right?

OpenStudy (anonymous):

subtract 2x from both sides and add all of -y'sin(x+y) to both sides. That will get everything y' on the right and everything else on the left.

OpenStudy (anonymous):

so we have -sin(x+y) -2x = -y'sin(x+y) +2y' ?

OpenStudy (anonymous):

Careful, its now a positive y'sin(x+y)

OpenStudy (anonymous):

true thanks!, and then can we divide both sides by sin (x+y)?

OpenStudy (freckles):

did you guys ever differentiate the 2x term?

OpenStudy (anonymous):

Ah, didnt even notice, lol. Good catch.

OpenStudy (anonymous):

Yeah, should be -2 on the left.

OpenStudy (anonymous):

so just to be clear what do you have so far as the equation?

OpenStudy (anonymous):

\[-\sin(x+y) -2 = y'\sin(x+y) + 2y'\]

OpenStudy (anonymous):

ok and can we divide both sides by sin(x+y)?

OpenStudy (anonymous):

Well, we want to factor out y' on the right hand side.

OpenStudy (anonymous):

so, y'(sin x+y+y') ?

OpenStudy (anonymous):

y'[sin(x+y) + 2] Do you see how I get that?

OpenStudy (anonymous):

yes I got it :)

OpenStudy (anonymous):

So then we would divide both sides by [sin(x+y) + 2]

OpenStudy (anonymous):

so, y' = -sin (x+y) -2/ [sin (x+y) +2 and can I cancel most of that out?

OpenStudy (anonymous):

Which would leave you with? :)

OpenStudy (anonymous):

y'= 0 ?

OpenStudy (anonymous):

Division can only equal 0 if the numerator is 0.

OpenStudy (anonymous):

-sin (x+y)/ sin (x+y)

OpenStudy (anonymous):

\[y' = \frac{ -\sin(x+y)-2 }{ \sin(x+y)+2}\] Your 2's are still there.

OpenStudy (anonymous):

\[y' =\frac{ -[\sin(x+y)+2] }{ \sin(x+y) + 2 } =?\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

And what would that cancel out to?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

oh ok -1

OpenStudy (anonymous):

So if y' = -1, do you know what that tells us?

OpenStudy (anonymous):

I'm not sure can we plug that number in to the original equation? not sure

OpenStudy (anonymous):

Well, we don't have to plug anything in. The idea is that a derivative gives us an equation for slope at any x-value. But we differentiated and got simply got -1. This means the slope of our original function is ALWAYS -1, no matter which point we choose. That's why we take the derivative, so we can figure out a slope. But now it wants us to construct a tangent line. Do you recall what we need to do to construct one?

OpenStudy (anonymous):

we needed our slope which we got, and we need to get the y intercept at (pi,pi)

OpenStudy (anonymous):

this way we can get our tangent line equation

OpenStudy (anonymous):

Well, a y-intercept would help us construct a line, but that doesn't always help us. We need a line through a "specific point", namely (pi,pi). Now since the slope is always -1, the point doesn't end up mattering as much, but we still always want to use the point we're given. The idea is we want to use the point-slope form formula, this will help us get a line through a specific point. So the point-slope equation is: \[y-y_{1} = m(x-x_{1})\] So just plug in the given point and you're slope into the equation and simplify :)

OpenStudy (anonymous):

I ended up with y = -x+2pi

OpenStudy (anonymous):

Exactly, that's your tangent line.

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (anonymous):

Yep, np :)